A ruler is 30 cm long.This is 25 cm less than 2/5 of the length of a large box.Shat is the length of the box?
Answer by stanbon(75887)

A ruler is 30 cm long.This is 25 cm less than 2/5 of the length of a large box.
What is the length of the box?
--------------------
Equation:
30 = (2/5)L -25
----
(2/5)L = 55
---
L = (5/2)55
L = 137.5 cm
-----------
Cheers,
Stan H.
----------